# Bounds and Bugs: The Limits of Symmetry Metrics to Detect Partisan Gerrymandering

### v3

@@ 2.1 Counting Districts Won
That the focus of a partisan gerrymander is to create an extreme number of districts won by one party is a point of agreement among quantitative social scientists, political scientists, mathematicians, and the general public. In their recent article [34], Katz, King, and Rosenblatt state that “Partisan gerrymanderers use their knowledge of voter preferences and their ability to draw favorable redistricting plans to maximize their party’s seat share.” They go on to emphasize that a gerrymanderer cares not about issues such as turnout or electoral responsiveness “unless it helps them win more seats.” In his “Defining a Gerrymander” section of [30], Grofman explains that gerrymandering results in one party winning an extreme number of seats compared to the number they might win in “a plan drawn on the basis of neutral principles.” In [49], Wang describes the effects of partisan gerrymandering on the US House in 2012 by the number of seats won by each party, and states that “In summary, the effects of partisan redistricting exceeded the amount of asymmetry caused by natural patterns of population.” And in [15], DeFord et all state that “the express intent of a partisan gerrymander is to secure for their party as many seats as possible under the constraints of voter geography and the other rules of redistricting.” While several of these authors caveat that what counts as “extreme” may depend on a region’s political geography [^3] We acknowledge that a state’s political geography impacts what specific numbers may be considered “extreme number of districts won.” This is why we conduct our empirical analysis in Section 5; in that section, we produce extreme maps using a state’s political geography., they all agree that number of districts won is the foundation of a partisan gerrymander.
@@
We also agree with Wang [49] and DeFord [15] that the political geography of a region being redistricted has significant impact on what number of districts won should be considered “extreme.” For example, Duchin et al showed that it was virtually impossible for Massachusetts to have a US House District with a majority of Republican voters, due to the political geography of that state [19]. Thus, a result of 9 out of 9 seats for Democrats should not be flagged as extreme at all for Massachusetts. We take into account each state’s political geography in our empirical analyses in Section 5.
@@ 3.1 The Seats-Votes Curve
The correspondence between the $MM$, $PB$, and the seats-votes curve described above is only true when the curve is drawn under the assumption of uniform partisan swing. Of course, uniform partisan swing is by no means guaranteed or even likely to happen in practice (see [52] for other models of swing). Grofman and King [29] do not even define the Partisan Bias as we do in Definition 2, and instead define it using the intersection of the seats-votes curve with the line $V=0.5$, where the seats-votes curve is constructed not by assuming uniform partisan swing, but by using statistical methods focused on the rank order [^5] Rank is in terms of partisan support for a particular political party. of districts. Grofman and King acknowledge that Partisan Bias is frequently defined as in our Definition 2, but call the definition of the Partisan Bias which uses the aforementioned statistical technique of defining the seats-votes curve the “state of the art” version of the metric [29]. However, as DeFord et al point out, complex definitions (such as Grofman and King’s “state of the art” version of Partisan Bias) must be “translated to the real-world ‘political thicket’ of redistricting” [16]. Wang agrees, by stating that a manageable standard to detect partisan gerrymandering “should be able to be clearly stated without case-specific or mathematics-intensive assumptions” [49].
@@ 4.1 Bounds: Partisan Bias
Suppose $(V,S)$ is a pair of rational numbers with $0\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Then there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Partisan Bias $PB$ for any $PB$ in the following ranges:
@@
Theorem 1 gives bounds on the value of the Partisan Bias when the number of districts is unlimited. However, using the same arguments as in the proofs in Appendix A, we have the following bounds for a fixed number of districts:
@@
Suppose $(V,S)$ is a pair of rational numbers with $0\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Suppose that $n$ is the number of districts and that $\ell$ is the number of districts lost by the party in consideration, so that $S=\frac{n-\ell}{n}$. Then there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Partisan Bias $PB$ for any $PB$ in the following ranges:
@@ 4.2 Bounds: Mean-Median Difference
Theorem 3 gives bounds on the value of the Mean-Median Difference when the number of districts is unlimited. However, using the same arguments as in the proofs in Appendix A, we have the following bounds for a fixed number of districts:
@@
Suppose $(V,S)$ is a pair of rational numbers with $0\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Suppose that $n$ is the number of districts and that $\ell$ is the number of districts lost by the party in consideration, so that $S=\frac{n-\ell}{n}$.
@@ 5 Bugs
In order to test whether the Mean-Median Difference and the Partisan Bias have extreme values on maps with an extreme number of districts won, we need to produce maps with an extreme number of districts won. We do this with a recently developed hill-climbing method called the “Short Burst Method,” developed by Cannon et al [10]. This hill-climbing process uses GerryChain, which is a Python library that implements a well-researched and widely accepted Markov Chain Monte Carlo (MCMC) process to create an ensemble of potential redistricting maps [18, 40]. In [10], Cannon et al showed that this hill-climbing method worked better than a biased random walk [^14] A biased random walk is the most common method of using a Markov Chain to find local or global maxima; in this case, we are searching for redistricting maps with a near-maximum number of districts won by some party. at finding maps with an extreme number of majority-minority districts. Our application is a direct parallel; while Cannon et al were looking for maps with an extreme number of districts which were majority-minority, here we are looking for maps with an extreme number of districts which are majority-Democrat or majority-Republican.
@@ 5.2 Results of Empirical Study
As mentioned in Section 5.1, we have results for both the Democratic party and the Republican Party for three different maps on each of six states. Note that, as explained in Section 5.1, the method we use to produce our maps is a hill-climbing method recently developed by Cannon et al [10] which utilizes the GerryChain library. Thus, the maps under consideration are not constructed in the same way as those constructed for a typical ensemble analysis (see Section 5.1 for details) and are not intended to provide representative samples from a specified distribution. Instead, they represent a large set of examples that have been optimized for extreme seats outcomes.
@@ 7 Acknowledgements
Appendix A Proofs of Theorems 1, 3, and 5
@@
Suppose $(V,S)$ is a pair of rational numbers with $S\geq\frac{1}{2}$, $S\leq 2V$, and $S\geq 2V-1$.
@@
Let $S=\frac{n-\ell}{n}$, and choose the denominator $n$ to be large. Choose $\epsilon>0$ small enough so that $\frac{1}{2}-(n-\ell)\epsilon>0$. Set $n-\ell$ districts to have vote share $V+\epsilon=\frac{1}{2}+\epsilon$; these are winning districts. Set one losing district to have vote share $V-(n-\ell)\epsilon=\frac{1}{2}-(n-\ell)\epsilon$, and the remaining losing districts to have vote share $V=\frac{1}{2}$. This constructed election has vote share $V$, seat share $S$, a maximum number of districts with vote share above $V$, and a minimum number of districts with vote share below $V$. The Partisan Bias is
$\frac{1}{2}\left(\frac{n-\ell}{n}-\frac{1}{n}\right)$
which has a supremal limit of $\frac{S}{2}$.
Now choose $\epsilon>0$ small enough so that $\frac{1}{2}+\ell\epsilon<1$. Set $\ell$ districts to have vote share $V-\epsilon=\frac{1}{2}-\epsilon$; these are losing districts. Set one losing district to have vote share $V+\ell\epsilon=\frac{1}{2}+\ell\epsilon$, and the remaining winning districts to have vote share $V=\frac{1}{2}$. This constructed election has vote share $V$, seat share $S$, a minimum number of districts with vote share above $V$, and a maximum number of districts with vote share below $V$. The Partisan Bias is
$\frac{1}{2}\left(\frac{1}{n}-\frac{\ell}{n}\right)$
which has an infimal limit of $\frac{S-1}{2}$. Thus, we have shown that for $V=\frac{1}{2}$, we can construct election data with PB for any
$\frac{S-1}{2}$ $<PB<\frac{S}{2}$
@@
Let $\ell$ be the number of losing districts so that $S=\frac{n-\ell}{n}$. Let $p$ be the number between 00 and 1 such that
$0\cdot p+V\left(1-p-\frac{n-\ell}{n}\right)+\frac{n-\ell}{n}\cdot\frac{1}{2}=V$
Note that this implies
$p=\frac{n-\ell}{n}\left(\frac{1}{2V}-1\right)=S\left(\frac{1}{2V}-1\right)$
By finding a common denominator, we can assume that that $p$’s denominator is $n$. Under the assumption that turnout in all districts is equal, having a proportion $p$ of districts having vote share 0, $S$ districts having vote share $\frac{1}{2}$ (considered winning districts), and $1-p-\frac{n-\ell}{n}$ districts having vote share $V$ gives the statewide vote share as $V$ and seat share $S$. It also maximizes the proportion of districts having vote share $V$.
@@
Thus, if we want to make PB as large as possible for an election winning $S$ districts and vote share $V$, we first set $n$ (the number of districts) to be very large. Then construct election data with a proportion $p$ of districts having vote share 0, $S$ districts having vote share $\frac{1}{2}$ (considered winning districts), nearly $1-p-\frac{n-\ell}{n}$ districts being $\epsilon$-larger than $V$, and a single district less than $V$ for an extremely small $\epsilon$.
If we want to make PB as small as possible for an election winning $S$ districts and vote share arbitrarily close to $V$, we first set $n$ (the number of districts) to be very large. Then construct election data with a proportion $p$ of districts having vote share 0, $S$ districts having vote share $\frac{1}{2}$ (considered winning districts), nearly $1-p-\frac{n-\ell}{n}$ districts being $\epsilon$-smaller than $V$, and a single district larger than $V$ for an extremely small $\epsilon$.
Using these two constructions, we have for $(V,S)$ the following infimal and supremal bounds:
$\frac{1}{2}\left(\frac{n-\ell}{n}-\left(1-p-\frac{n-\ell}{n}+p\right)\right)$ $<PB<\frac{1}{2}\left(\frac{n-\ell}{n}+1-p-\frac{n-\ell}{n}-p\right)$
$\frac{1}{2}\left(2S-1\right)$ $<PB<\frac{1}{2}\left(1-2p\right)=\frac{1}{2}\left(1-2S\left(\frac{1}{2V}-1\right)\right)$
$S-\frac{1}{2}$ $<PB<\frac{1}{2}-S\left(\frac{1}{2V}-1\right)$
Case 3: $\frac{1}{2}<V\leq 1$
Let $\ell$ be the number of losing districts so that $S=\frac{n-\ell}{n}$. Let $p$ be the number between 00 and 1 such that
$\frac{\ell}{n}\cdot\frac{1}{2}+V\left(1-p-\frac{\ell}{n}\right)+p=V$
Note that this implies
$p=\frac{\frac{\ell}{n}\left(V-\frac{1}{2}\right)}{1-V}=\frac{(1-S)\left(V-\frac{1}{2}\right)}{1-V}$
By finding a common denominator, we can assume that that $p$’s denominator is $n$. Under the assumption that turnout in all districts is equal, having a proportion $p$ of districts having vote share 1, $\frac{\ell}{n}=1-S$ districts having vote share $\frac{1}{2}$ (considered losing districts), and $1-p-\frac{\ell}{n}$ districts having vote share $V$ gives the statewide vote share as $V$ and seat share $S$. It also maximizes the proportion of districts having vote share $V$.
Thus, to make PB as large as possible for an election winning $S$ districts and vote share $V$, we would first set $n$ (the number of districts) to be very large. Then construct election data with a proportion $p$ of districts having vote share 1, $\frac{\ell}{n}=1-S$ districts having vote share $\frac{1}{2}$ (considered losing districts), nearly $1-p-\frac{\ell}{n}$ districts being $\epsilon$-larger than $V$, and a single district smaller than $V$ for an extremely small $\epsilon$
If we want to make PB as small as possible for an election winning $S$ districts and vote share arbitrarily close to $V$, we would first set $n$ (the number of districts) to be very large. Then construct election data with a proportion $p$ of districts having vote share 1, $\frac{\ell}{n}=1-S$ districts having vote share $\frac{1}{2}$ (considered losing districts), nearly $1-p-\frac{\ell}{n}$ districts being $\epsilon$-smaller than $V$, and a single district larger than $V$ for an extremely small $\epsilon$.
Using these two constructions, we have for $(V,S)$ the following infimal and supremal bounds:
$\frac{1}{2}\left(p-\left(1-p-\frac{\ell}{n}+\frac{\ell}{n}\right)\right)$ $<PB<\frac{1}{2}\left(p+1-p-\frac{\ell}{n}-\frac{\ell}{n}\right)$
$\frac{1}{2}\left(2p-1\right)$ $<PB<\frac{1}{2}\left(1-2\frac{\ell}{n}\right)$
$\frac{(1-S)\left(V-\frac{1}{2}\right)}{1-V}-\frac{1}{2}$ $<PB<S-\frac{1}{2}$
@@
Since $S=\frac{1}{2}$, the smallest that the median value could be is $\frac{1}{4}$, which would occur when the middle vote share values are 0 (a losing district) and $\frac{1}{2}$ (a winning district). For $\frac{1}{4}\leq V<\frac{1}{2}$, this median value is achievable by letting the other winning vote shares range from $\frac{1}{2}$ (when $V=\frac{1}{4}$) to 1 (for $V$ approaching $\frac{1}{2}$). This gives a minimum MM of $\frac{1}{4}-V$.
For $\frac{1}{2}\leq V\leq\frac{3}{4}$, let $p$ be a value such that
@@
$p=2\left(V+\frac{1}{2n}-\frac{1}{2}\right)\approx 2V-1$
Then the smallest that the median value can be for a constructed election with seat share $S=\frac{1}{2}$ and vote share $V$ is a value approaching
$\frac{p+\frac{1}{2}}{2}=\frac{2V-1+\frac{1}{2}}{2}=V-\frac{1}{4}$
Thus, the infimal Mean-Median difference for $\frac{1}{2}\leq V\leq\frac{3}{4}$ is $-\frac{1}{4}$. Using symmetry, we get
@@
Case 2: $S>\frac{1}{2}$, $\frac{1}{4}\leq V<\frac{3}{4}$
@@
Firstly note that, since we assume that $S>\frac{1}{2}$, the median vote share must be at least as large as $\frac{1}{2}$. For $V\leq\frac{1}{2}$, we can achieve a median vote share of $\frac{1}{2}$ by constructing $S$ of the districts to have vote share $\frac{1}{2}$ (considered winning districts) and choosing the losing districts to have vote shares so that $\frac{1}{n}\sum_{i=1}^{n}V_{i}=V$, since we must have $S\leq 2V$. This would give vote share $V$, seat share $S$, and the smallest possible median of $\frac{1}{2}$.
For $\frac{1}{2}<V<\frac{3}{4}$, we can choose half of the districts to have vote share $\frac{1}{2}$ ($\ell$ of those considered to be losing districts, the rest winning) and then half to have vote share $V+\left(V-\frac{1}{2}\right)$. This would give vote share $V$, seat share $S$, and the smallest possible median of $\frac{1}{2}$.

### v4

@@ 2.1 Counting Districts Won
That the focus of a partisan gerrymander is to create an extreme number of districts won by one party is a point of agreement among quantitative social scientists, political scientists, mathematicians, and the general public. In their recent article [34], Katz, King, and Rosenblatt state that “Partisan gerrymanderers use their knowledge of voter preferences and their ability to draw favorable redistricting plans to maximize their party’s seat share.” They go on to emphasize that a gerrymanderer cares not about issues such as turnout or electoral responsiveness “unless it helps them win more seats.” In his “Defining a Gerrymander” section of [30], Grofman explains that gerrymandering results in one party winning an extreme number of seats compared to the number they might win in “a plan drawn on the basis of neutral principles.” In [49], Wang describes the effects of partisan gerrymandering on the US House in 2012 by the number of seats won by each party, and states that “In summary, the effects of partisan redistricting exceeded the amount of asymmetry caused by natural patterns of population.” And in [15], DeFord et al. state that “the express intent of a partisan gerrymander is to secure for their party as many seats as possible under the constraints of voter geography and the other rules of redistricting.” While several of these authors caveat that what counts as “extreme” may depend on a region’s political geography [^3] We acknowledge that a state’s political geography impacts what specific numbers may be considered “extreme number of districts won.” This is why we conduct our empirical analysis in Section 5; in that section, we produce extreme maps using a state’s political geography., they all agree that number of districts won is the foundation of a partisan gerrymander.
@@
We also agree with Wang [49] and DeFord [15] that the political geography of a region being redistricted has significant impact on what number of districts won should be considered “extreme.” For example, Duchin et al. showed that it was virtually impossible for Massachusetts to have a US House District with a majority of Republican voters, due to the political geography of that state [19]. Thus, a result of 9 out of 9 seats for Democrats should not be flagged as extreme at all for Massachusetts. We take into account each state’s political geography in our empirical analyses in Section 5.
@@ 3.1 The Seats-Votes Curve
The correspondence between the $MM$, $PB$, and the seats-votes curve described above is only true when the curve is drawn under the assumption of uniform partisan swing. Of course, uniform partisan swing is by no means guaranteed or even likely to happen in practice (see [52] for other models of swing). Grofman and King [29] do not even define the Partisan Bias as we do in Definition 2, and instead define it using the intersection of the seats-votes curve with the line $V=0.5$, where the seats-votes curve is constructed not by assuming uniform partisan swing, but by using statistical methods focused on the rank order [^5] Rank is in terms of partisan support for a particular political party. of districts. Grofman and King acknowledge that Partisan Bias is frequently defined as in our Definition 2, but call the definition of the Partisan Bias which uses the aforementioned statistical technique of defining the seats-votes curve the “state of the art” version of the metric [29]. However, as DeFord et al. point out, complex definitions (such as Grofman and King’s “state of the art” version of Partisan Bias) must be “translated to the real-world ‘political thicket’ of redistricting” [16]. Wang agrees, by stating that a manageable standard to detect partisan gerrymandering “should be able to be clearly stated without case-specific or mathematics-intensive assumptions” [49].
@@ 4.1 Bounds: Partisan Bias
Suppose $(V,S)$ is a pair of rational numbers with $\frac{1}{4}\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$.
If $S=1$ and either $V=\frac{1}{2}$ or $V=1$, then we must have $PB=0$.
Otherwise, there exists constructed election data with vote share $V$, seat share $S$, turnout is equal in all districts, and Partisan Bias $PB$ for any $PB$ in the following ranges:
@@
Theorem 1 gives bounds on the value of the Partisan Bias when the number of districts is unlimited. We have the following bounds for a fixed number of districts:
@@
Suppose $(V,S)$ is a pair of rational numbers with $\frac{1}{4}\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Suppose that $n$ is the number of districts; that is, $S$ can be written as a rational number with denominator $n$.
If $S=1$ and either $V=\frac{1}{2}$ or $V=1$, then we must have $PB=0$.
Otherwise, there exists constructed election data with vote share $V$, seat share $S$, $n$ districts, turnout is equal in all districts, and Partisan Bias $PB$ for any $PB$ in the following ranges:
[…]
@@ 4.2 Bounds: Mean-Median Difference
Theorem 3 gives bounds on the value of the Mean-Median Difference when the number of districts is unlimited. We have the following bounds for a fixed number of districts:
@@
Suppose $(V,S)$ is a pair of rational numbers with $\frac{1}{4}\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Suppose that $n\geq 3$ is the number of districts and that $\ell$ is the number of districts lost by the party in consideration, so that $S=\frac{n-\ell}{n}$.
@@ 5 Bugs
In order to test whether the Mean-Median Difference and the Partisan Bias have extreme values on maps with an extreme number of districts won, we need to produce maps with an extreme number of districts won. We do this with a recently developed hill-climbing method called the “Short Burst Method,” developed by Cannon et al. [10]. This hill-climbing process uses GerryChain, which is a Python library that implements a well-researched and widely accepted Markov Chain Monte Carlo (MCMC) process to create an ensemble of potential redistricting maps [18, 40]. In [10], Cannon et al. showed that this hill-climbing method worked better than a biased random walk [^14] A biased random walk is the most common method of using a Markov Chain to find local or global maxima; in this case, we are searching for redistricting maps with a near-maximum number of districts won by some party. at finding maps with an extreme number of majority-minority districts. Our application is a direct parallel; while Cannon et al. were looking for maps with an extreme number of districts which were majority-minority, here we are looking for maps with an extreme number of districts which are majority-Democrat or majority-Republican.
@@ 5.2 Results of Empirical Study
As mentioned in Section 5.1, we have results for both the Democratic party and the Republican Party for three different maps on each of six states. Note that, as explained in Section 5.1, the method we use to produce our maps is a hill-climbing method recently developed by Cannon et al. [10] which utilizes the GerryChain library. Thus, the maps under consideration are not constructed in the same way as those constructed for a typical ensemble analysis (see Section 5.1 for details) and are not intended to provide representative samples from a specified distribution. Instead, they represent a large set of examples that have been optimized for extreme seats outcomes.
@@ 7 Acknowledgements
Appendix A Proofs
For ease of exposition, we use the notation that a district has vote share $\frac{1}{2}^{+}$ if party $A$’s vote share is $\frac{1}{2}$ and party $A$ wins that district. Similarly, a district has vote share $\frac{1}{2}^{-}$ if party $A$’s vote share is $\frac{1}{2}$ and party $A$ loses that district.
@@ Appendix A Proofs
Suppose $(V,S)$ is a pair of rational numbers with $\frac{1}{4}\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Let $n$ be the number of districts; that is, $S$ can be written as a rational number with denominator $n$.
@@
Suppose first that $S=1$. In this case, all districts must have vote share $\frac{1}{2}^{+}$, which means that the PB of this election is 0.
Otherwise, say $S=\frac{n-\ell}{n}$ for $\ell\geq 1$. In this case, choose $\epsilon>0$, $\epsilon\in{\mathbb{Q}}$ small enough so that $\frac{1}{2}-(n-\ell)\epsilon>0$. Set $n-\ell$ districts to have vote share $V+\epsilon=\frac{1}{2}+\epsilon$; these are winning districts. Set one losing district to have vote share $V-(n-\ell)\epsilon=\frac{1}{2}-(n-\ell)\epsilon$, and the remaining losing districts to have vote share $V=\frac{1}{2}^{-}$. This constructed election has vote share $V$, seat share $S$, a maximum number of districts with vote share above $V$, and a minimum number of districts with vote share below $V$. The Partisan Bias is
$\frac{1}{2}\left(\frac{n-\ell}{n}-\frac{1}{n}\right)=\frac{1}{2}\left(S-\frac{1}{n}\right)$
as claimed in the Theorem statement.
Now choose $\epsilon>0$, $\epsilon\in{\mathbb{Q}}$ small enough so that $\frac{1}{2}+\ell\epsilon<1$. Set $\ell$ districts to have vote share $V-\epsilon=\frac{1}{2}-\epsilon$; these are losing districts. Set one winning district to have vote share $V+\ell\epsilon=\frac{1}{2}+\ell\epsilon$, and the remaining winning districts to have vote share $V=\frac{1}{2}^{+}$. This constructed election has vote share $V$, seat share $S$, a minimum number of districts with vote share above $V$, and a maximum number of districts with vote share below $V$. The Partisan Bias is
$\frac{1}{2}\left(\frac{1}{n}-\frac{\ell}{n}\right)=\frac{1}{2}\left(S-1+\frac{1}{n}\right)$
as claimed in the Theorem statement.
@@
Let $k=\lceil n\left(1-\frac{S}{2V}\right)-1\rceil$. Note that $k$ is chosen as the largest integer such that
$V\cdot\frac{k}{n}+S\cdot\frac{1}{2}<V$
To maximize PB on an election with seat share $S$ and vote share $V$, set proportion $S$ districts to have vote share $\frac{1}{2}+\delta$, and set $k$ districts to have vote share $V+\epsilon$ for $\delta,\epsilon\in{\mathbb{Q}}$, $\delta,\epsilon>0$ small enough that $(V+\epsilon)\cdot\frac{k}{n}+S\cdot\left(\frac{1}{2}+\delta\right)=V$ and $V+\epsilon\leq\frac{1}{2}$, $\frac{1}{2}+\delta\leq 1$. (Such $\epsilon$ and $\delta$ are guaranteed by the choice of $k$, along with the fact that $V<\frac{1}{2}\leq S$.) The remaining districts will have vote share 0. This maximal PB value is $\frac{1}{2}\left(S+\frac{k}{n}-\left(1-S-\frac{k}{n}\right)\right)=\frac{1}{2}\left(2S-1+\frac{2k}{n}\right)$.
If we want to make PB as small as possible on an election with seat share $S$ and vote share $V$, we recall that $S\leq 2V$, so that the $p\in{\mathbb{Q}}$ satisfying
$p(1-S)+S\cdot\frac{1}{2}=V$
must have $p\geq 0$. Since $V<\frac{1}{2}$, we must have $p<V$. Thus, the minimal PB achievable is $\frac{1}{2}\left(S-(1-S)\right)=S-\frac{1}{2}$.
Using these two constructions, we have for $(V,S)$ the following bounds:
$S-\frac{1}{2}$ $\leq PB\leq\frac{1}{2}\left(2S-1+\frac{2k}{n}\right)$
$S-\frac{1}{2}$ $\leq PB\leq\frac{1}{2}\left(2S-1+\frac{2\left\lceil n\left(1-\frac{S}{2V}\right)-1\right\rceil}{n}\right)$
Case 3: $\frac{1}{2}<V<1$, $S=1$
In this case, we can choose $\epsilon>0,\epsilon\in{\mathbb{Q}}$ small enough that $V+(n-1)\epsilon\leq 1$ and $V-(n-1)\epsilon\geq\frac{1}{2}$. The largest PB is achieved with $n-1$ districts having vote share $V+\epsilon$ and one district having vote share $V-(n-1)\epsilon$, resulting in a PB value of $\frac{1}{2}\left(\frac{n-1}{n}-\frac{1}{n}\right)=\frac{n-2}{2n}$. And the smallest PB is achieved with $n-1$ districts having vote share $V-\epsilon$ and one district having vote share $V+(n-1)\epsilon$, resulting in a PB value of $\frac{1}{2}\left(\frac{1}{n}-\frac{n-1}{n}\right)=\frac{2-n}{2n}$.
Case 4:. $\frac{1}{2}<V<1$, $S<1$
@@
Here, if we want to make PB as large as possible for an election winning seat share $S$ and vote share $V$, we construct election data with proportion $S$ districts having vote share $V+\epsilon$ for $\epsilon\in{\mathbb{Q}},\epsilon>0$, and proportion $1-S$ districts having vote share $\frac{1}{2}-\delta$ for $\delta\in{\mathbb{Q}},\delta\geq 0$ so that $V+\epsilon\leq 1$, $\frac{1}{2}-\delta\geq 0$, and
$\left(1-S\right)\left(\frac{1}{2}-\delta\right)+S\left(V+\epsilon\right)=V.$
Such $\epsilon$ and $\delta$ are guaranteed by the facts that $S<1$ and $S\geq 2V-1$. This results in a maximal PB of $\frac{1}{2}\left(S-(1-S)\right)=S-\frac{1}{2}$.
If we want to make PB as small as possible for an election winning $S$ districts and vote share $V$, we set $k=\lceil n\left(1-\frac{1-S}{2(1-V)}\right)-1\rceil$. Note that $k$ is chosen as the largest integer such that
$(1-S)\cdot\frac{1}{2}+\frac{k}{n}V+\left(S-\frac{k}{n}\right)>V$
Choose $\epsilon\in{\mathbb{Q}},\delta\in{\mathbb{Q}}$, and $\gamma\in{\mathbb{Q}}$ with $\epsilon>0,\delta,\gamma\geq 0$ so that $(1-S)\cdot\left(\frac{1}{2}-\delta\right)+\frac{k}{n}(V-\epsilon)+\left(S-\frac{k}{n}\right)\left(1-\gamma\right)=V$, $\frac{1}{2}-\delta\geq 0$, $V-\epsilon\geq\frac{1}{2}$ and $1-\gamma>V$. Such $\epsilon,\delta$, and $\gamma$ exist because of the choice of $k$ and the facts that $V>\frac{1}{2}$ and $S<1$. Thus, to minimize PB, we construct election data with a proportion $S-\frac{k}{n}$ of districts having vote share $1-\gamma$, a proportion of $1-S$ districts having vote share $\frac{1}{2}-\delta$, and $k$ districts having vote share $V-\epsilon$. This results in a PB value of $\frac{1}{2}\left(S-\frac{k}{n}-\left(1-S+\frac{k}{n}\right)\right)=\frac{1}{2}\left(2S-1-\frac{2k}{n}\right)$.
Using these two constructions, we have for $(V,S)$ the following bounds:
$\frac{1}{2}\left(2S-1-\frac{2k}{n}\right)$ $\leq PB\leq S-\frac{1}{2}$
$\frac{1}{2}\left(2S-1-\frac{2\left\lceil n\left(1-\frac{1-S}{2(1-V)}\right)-1\right\rceil}{n}\right)$ $\leq PB\leq S-\frac{1}{2}$
Case 5: $V=1$. In this case, $S=1$ and all districts have vote share 1. Thus, the only value for PB is 1.
@@
Take the limit as $n\to\infty$ in Theorem 2.
Proof of Theorem 4.
Suppose $(V,S)$ is a pair of rational numbers with $\frac{1}{4}\leq V\leq 1$, $\frac{1}{2}\leq S\leq 1$, $S\leq 2V$, and $S\geq 2V-1$. Suppose that $n\geq 3$ is the number of districts and that $\ell$ is the number of losing districts, so that $S=\frac{n-\ell}{n}$.
@@
Since $S=\frac{1}{2}$, the smallest that the median value could be is $\frac{1}{4}$, which would occur when the middle vote share values are 0 (a losing district) and $\frac{1}{2}^{+}$ (a winning district). For $V\geq\frac{1}{4}$, this median value is achievable by letting the other winning vote shares range from $\frac{1}{2}$ (when $V=\frac{1}{4}$) to 1 (for $V$ approaching $\frac{1}{2}$). This gives a minimum MM of $\frac{1}{4}-V$, which is achievable for $V\geq\frac{1}{4}$, so long as $V\leq\frac{0+\frac{1}{2}+\frac{n}{2}-1}{n}=\frac{n-1}{2n}$.
For $\frac{n-1}{2n}\leq V\leq\frac{3}{4}$, let $p\in{\mathbb{Q}}$ be the value such that
@@
$p=2\left(V+\frac{1}{2n}-\frac{1}{2}\right)$
Then for $\frac{n-1}{2n}\leq V\leq\frac{3}{4}$ the smallest that the median value can be for a constructed election with seat share $S=\frac{1}{2}$ and vote share $V$ occurs when there are $\frac{n}{2}-1$ districts with vote share 1, 1 district with vote share $\frac{1}{2}^{+}$, and $\frac{n}{2}$ districts with vote share $p$. This results in a MM value of
$\frac{p+\frac{1}{2}}{2}-V=\frac{2\left(V+\frac{1}{2n}-\frac{1}{2}\right)+\frac{1}{2}}{2}-V=\frac{1}{2n}-\frac{1}{4}.$
Using symmetry, we get
@@
[…]
Case 2: $S>\frac{1}{2}$, $\frac{1}{4}<V\leq 1$, lower bounds for MM
@@
Firstly note that, since we assume that $S>\frac{1}{2}$, the median vote share must be at least as large as $\frac{1}{2}$. For $V\leq\frac{1}{2}$, we can achieve a median vote share of $\frac{1}{2}$ by constructing $S$ of the districts to have vote share $\frac{1}{2}^{+}$ and choosing the losing districts to have the vote share $p\in{\mathbb{Q}}$ satisfying $p(1-S)+\frac{1}{2}\cdot S=V$. This $p$ will have $p\leq V\leq\frac{1}{2}$, since $V\leq\frac{1}{2}$. This would give vote share $V$, seat share $S$, and the smallest possible median of $\frac{1}{2}$.
For $\frac{1}{2}<V$, first suppose that $n$ is even (so that $n\geq 4$). Then we can choose $\ell$ districts to have vote share $\frac{1}{2}^{-}$ and $\frac{n}{2}+1-\ell$ districts to have vote share $\frac{1}{2}^{+}$, giving the smallest possible median value of $\frac{1}{2}$. The remaining districts have vote share $p\in{\mathbb{Q}}$ which results in a statewide vote share of $V$. However this is only possible if the equation
$\frac{\left(\frac{n}{2}+1\right)\cdot\frac{1}{2}+p\left(\frac{n}{2}-1\right)}{n}=V$